OpenStudy (anonymous):

Suppose the equation, y = 21(1.06)x, represents book sales in billions of dollars since 2000. X represents the number of years after 2000. The table below shows the book sales for the four years after 2000, to the nearest billion. Which year has an incorrect amount? Years after 2000 (x) 1 2 3 4 Billions of dollars (y) 23 24 25 27

OpenStudy (anonymous):

A. 1 23 B. 2 24 C. 3 25 D. 4 27

OpenStudy (anonymous):

All you have to do is plug in the values of the equation into a calculator

OpenStudy (anonymous):

@yazthespaz for my answer U got B. 24 But I don't know if it right?

OpenStudy (anonymous):

Oh okay i will check the answer for you

OpenStudy (anonymous):

@yazthespaz thanks I did it on my sheet.

OpenStudy (anonymous):

It's actually A and i will explain why now

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

y = 21(1.06)x plug in A 23 = 21(1.06)1 23 = 21(1.06) 23 = 22.26 You round 22.26 to 23

OpenStudy (anonymous):

Oh I understand now grazie. (:

OpenStudy (anonymous):

No problem :)